# Local stability guarantees for data-driven quadratically nonlinear models

### v1

@@ I Introduction
To determine if a Galerkin model is long-term bounded, Schlegel and Noack [1] developed a “trapping theorem” with necessary and sufficient conditions for long-term model stability for systems that exhibit quadratic, energy-preserving nonlinearities. This theorem can be used as an effective diagnostic for evaluating the stability of reduced-order models. Moreover, the theorem was subsequently incorporated into system identification methods, for producing data-driven models that are provably long-term bounded [13]. The ability to promote global stability guarantees has broad potential for data-driven models, and has already been extended to neural-network-based system identification methods [14]. Recently, Goyal et al[15] extended the trapping theorem in quadratic systems to generalized energy-preserving nonlinearities, achieved through constraints of a special parametrization that guarantee the stability of the learned model by construction. For general quadratic systems, Kramer [16] discusses the condition for such systems to be locally stable by means of a Lyapunov function and provide better estimates of the radii of the domain of attraction by a novel optimization-based approach.
@@
In this work, we develop a theorem in the context of projection-based ROMs that enables us to promote local stability of data-driven models with quadratic nonlinearities. The theorem provides sufficient condition under which linear-quadratic systems with a relaxed quadratically energy-preserving structure are locally stable. We can also reproduce the Schlegel and Noack theorem [1, 13] by tightening the relaxation of the energy-preserving constraints. Formalizing stability under weak breaking of the energy-preserving nonlinearity is critical because many flows exhibit boundary conditions such as inflow/outflow conditions. Providing conditions under which this class of systems is guaranteed to be stable is helpful for designing fluid flow control.
@@ II Reduced-order modeling
The density $\rho$ and viscosity $\nu$ are assumed constant, and the pressure is denoted $p$. These equations are expected to accurately approximate incompressible flow of Newtonian fluids with constant density, kinematic viscosity, and inflow/outflow boundary conditions. Other complex fluid flows such as ocean currents, viscoelastic fluids, or ferrofluids are described by models that produce higher order polynomial nonlinearities or trigonometric nonlinear terms, but may be amenable to the quadratic analysis here if regimes exist where higher order nonlinearities are inactive or weak [26].
The goal of a projection-based ROM is to transform this high-dimensional system into a lower-dimensional system of size $r\ll n$ that captures the essential dynamics. One way to reduce the set of governing equations to a set of ordinary differential equations is by decomposition into a desired low-dimensional basis $\{\bm{\chi}_{i}(\bm{x})\}$ in a process commonly referred to as Galerkin expansion:
@@
The sum over $i$ is implied and $i\in\{1,...,r\}$. Here, $\overline{\bm{u}}(\bm{x})$ is a base flow which fulfills stationary boundary conditions, $\bm{\chi}_{i}(\bm{x})$ are spatial modes of the expansion for the fluctuation $\bm{u}^{\prime}(\bm{x},t)=\bm{u}(\bm{x},t)-\overline{\bm{u}}(\bm{x})$, and $a_{i}(t)$ describe how the amplitude of these modes vary in time. The proper orthogonal decomposition (POD) [10, 27] is frequently used to obtain the basis, since the modes $\bm{\chi}_{i}(\bm{x})$ are orthogonal and ordered by maximal energy content. The governing equations are then Galerkin projected onto the basis $\{\bm{\chi}_{i}(\bm{x})\}$ by substituting Eq. (2) into the PDE of Eq. (II) and using inner products with $\bm{\chi}_{k}$ to remove the spatial dependence. Orthogonal projection onto POD modes is the simplest and most common procedure, resulting in POD-Galerkin models, although Petrov-Galerkin projection [4, 5] improves model performance in some cases.
@@ II.1 Stability definitions
We begin with the definition of boundedness. In this paper, “boundedness” and “bounded” always imply the long-term behavior of the system. Additionally, we will use the terms “global stability” and “globally long-term bounded” interchangeably in this paper for convenience.
@@
$\dot{\bm{x}}=f(\bm{x})$ (3)
where map $f:\mathcal{D}\rightarrow\mathbb{R}^{n}$ is locally Lipschitz from $\mathcal{D}\subset\mathbb{R}^{n}$ onto $\mathbb{R}^{n}$.
@@ 1. $\bullet$
    (uniformly) bounded if there exists a constant $\rho>0$ that is independent of the initial time $t_{0}$, and for every $c\in(0,\rho)$, there is $\mu=\mu(c)>0$ such that
    $\|{\bm{x}}(t_{0})\|\leq c\ \Rightarrow\ \|{\bm{x}}(t)\|\leq\mu,\ \forall t\geq t_{0},$ (4)
@@ 2. $\bullet$
    globally (uniformly) bounded if (4) holds for arbitrarily large $\rho$.
While local or global boundedness is very general, we will show that for the case of quadratically nonlinear systems, boundedness can be addressed by considering a more specialized class of stability via trapping regions. Now we give the definition of a trapping region.
@@
Consider the dynamical system of (3) and let $\phi_{t}(\bm{x})$ denote a position of a solution of the system at time $t$ that started at $\bm{x}\in\mathcal{D}$, or a subsequent state at time $t$ of a trajectory. Then a set $\mathcal{N}\subset\mathcal{D}$ is
@@
    an attracting trapping region if there exists a compact set $E\subset\mathcal{D}$ and time $T$ such that $\mathcal{N}\subset E$ is a trapping region and
    $\phi_{t}(E\setminus\mathcal{N})\subset\mathrm{int}(\mathcal{N}),\ \forall t>T,$ (6)
@@ 3. $\bullet$
Figure 1: Sketch of an attracting trapping region. $\bm{\Gamma}_{1}$ (in red, solid) and $\bm{\Gamma}_{2}$ (in blue, solid) are two trajectories starting outside an attracting trapping region $\bm{\mathcal{N}}$ where all trajectories will eventually enter in and stay there forever. $V$ is a Lyapunov function and each dashed closed surface $V=V(\bm{x})=c_{i}\ (i=1,2,3,4)$ is a Lyapunov surface with $c_{1}>c_{2}>c_{3}>c_{4}$. For each Lyapunov surface outside $\bm{\mathcal{N}}$, $\bm{\Gamma}_{2}$ only passes through at most once, while $\bm{\Gamma}_{1}$ crosses $V=c_{3}$ more than once before each enters into $\mathcal{\bm{N}}$. If we consider the Lyapunov function as a form of energy, along $\bm{\Gamma}_{2}$, the energy is always decreasing before entering $\mathcal{\bm{N}}$. If all trajectories starting outside $\bm{\mathcal{N}}$ behave the same as $\bm{\Gamma}_{2}$, then $\bm{\mathcal{N}}$ is a monotonically attracting trapping region, otherwise an attracting trapping region.
@@ II.2 Flows with energy-preserving nonlinearities
$E_{i}$, $L_{ij}$, and $Q_{ijk}$ are tensors of time-independent coefficients, obtained from spatial inner products between the $\bm{\chi}_{i}(\bm{x})$ and the spatiotemporal operators that define the model dynamics. We have assumed that the pressure gradient term in Eq. (II) has been removed as in Rowley et al[28]. The class of systems considered for the original trapping theorem are those with energy-preserving nonlinearity, which is defined as
@@
$Q_{ijk}$ are symmetric in swapping $j$ and $k$ without loss of generality. To see why this condition preserves the energy, consider $\overline{\bm{u}}(\bm{x})$ being the mean flow and we can introduce the standard kinetic energy $K$,
@@ II.3 Schlegel and Noack trapping theorem
The Schlegel and Noack trapping theorem provides necessary and sufficient conditions for energy-preserving, effectively nonlinear, quadratic systems to exhibit a trapping region with upper bound size that can be estimated by $B(\bm{m},R_{m})$, a ball centered at $\bm{m}\in\mathbb{R}^{r}$ with radius $R_{m}$. Outside this region the rate of change of energy $K$ is negative everywhere, producing a Lyapunov function that renders this system globally stable. The trapping theorem begins by recentering the origin by a now arbitrary constant vector $\bm{m}$ so that we can define a shifted energy expressed in terms of the shifted state vector $\bm{y}(t)=\bm{a}(t)-\bm{m}$,
@@
Consider an effectively quadratically nonlinear system. There exists a monotonically trapping region at least as small as the ball $B(\bm{m},R_{m})$ if and only if the real, symmetric matrix $\bm{A}^{S}$ is negative definite (Hurwitz) with eigenvalues $\lambda_{r}\leq\cdots\leq\lambda_{1}<0$; the radius is then given by $R_{m}=\|\bm{d}\|_{2}/|\lambda_{1}|$. [^1] If a system is long-term bounded (not necessarily exhibiting a monotonically trapping region) and effectively nonlinear, the existence of an $\bm{m}$ producing a negative semidefinite $\bm{A}^{S}$ is still guaranteed.
In practice, the goal is then to find an origin $\bm{m}$ so that the matrix $\bm{A}^{S}$ is Hurwitz, guaranteeing a trapping region and global stability. There is a caveat for systems that do not exhibit effective nonlinearity; effective nonlinearity means that there is no linear subspace that a trajectory starts and remains in indefinitely, so that the quadratic nonlinearities $Q_{ijk}y_{j}y_{k}$ are never activated.
@@ II.4 Analytical estimates of the local stability of linear-quadratic systems
In this section, we use Lyapunov’s direct method to explore the stability of a linear-quadratic system of the form of Eq. (8). We consider the situation where the nonlinearities are no longer exactly energy-preserving for a linear-quadratic system. However, we assume that the breaking of the energy-preserving nonlinearity is bounded by a small amount $\epsilon_{Q}$, i.e.,
$\left|Q_{ijk}+Q_{jik}+Q_{kji}\right|\leq\epsilon_{Q},$ (17)
@@
where $\otimes$ denotes the standard Kronecker product. Notice that we previously assume that $Q_{ijk}$ are symmetric in the last two indices, hence $\bm{H}$ is symmetric, i.e., $\bm{H}(\bm{y}_{1}\otimes\bm{y}_{2})=\bm{H}(\bm{y}_{2}\otimes\bm{y}_{1})$ for arbitrary $\bm{y}_{1},\bm{y}_{2}\in\mathbb{R}^{r}$, which does not change the dynamics [30]. To generalize the discussion beyond the shifted energy, we define a Lyapunov function $v(\bm{y})$ given a Lyapunov matrix $\bm{P}$,
$v(\bm{y})$ $=\frac{1}{2}\langle\bm{y},\bm{y}\rangle_{\bm{P}}=\frac{1}{2}\bm{y}^{T}\bm{P}\bm{y},$ (23a)
@@
$\dot{v}(\bm{y})$ $=\bm{y}^{T}\bm{P}\dot{\bm{y}}.$ (24a)
Importantly, the traditional trapping theorem uses $\bm{P}=\bm{I}$, the identity, but for two-dimensional fluid flows the enstrophy is an invariant with $\bm{P}\neq\bm{I}$ and it can also be immediately used in the following results. Similar invariants exist in MHD, and more general Lyapunov matrices $\bm{P}$ may exist for certain flows. Section IV.3 will illustrate provably locally-stable, data-driven models of the von Karman vortex street using the enstrophy version of the various trapping theorems described in the present work.
Since the domain of boundedness $\bm{\Omega}_{\bm{m}}$ can be complicated geometrically, we consider closed balls $\textbf{B}\subseteq\bm{\Omega}_{\bm{m}}$ as the candidates of regions where the solutions of Eq. (22) are bounded,
$\textbf{B}_{\bm{y}}(\bm{R})=\{\bm{y}\in\mathbb{R}^{r}\ |\ \|\bm{y}\|\leq\bm{R}\}.$
We now provide a lower-bound for the domain of boundedness by the set $\textbf{B}_{\bm{y}}$. The task then becomes to make this bound as tight as possible, which is, to find $\sup{\bm{R}}$ as the estimate of the local stability radius. This leads to the following theorem.
@@
    $\bm{\tilde{A}}^{S}=\bm{P}\bm{A}^{S}$ is Hurwitz, i.e.
    $\lambda_{r}(\bm{\tilde{A}}^{S})\leq...\leq\lambda_{1}(\bm{\tilde{A}}^{S})<0$.
@@
    $\lambda^{2}_{1}-\frac{4r^{\frac{3}{2}}\epsilon_{Q}}{3}\|\bm{\tilde{d}}\|_{2}\geq 0$, where $\epsilon_{Q}=\|\tilde{Q}_{ijk}+\tilde{Q}_{ijk}+\tilde{Q}_{kji}\|_{\mathrm{max}}$, $\bm{\tilde{Q}}=\bm{PQ}$ and $\tilde{\bm{d}}=\bm{P}\bm{d}$.
@@
$\rho_{-}=\frac{3|\lambda_{1}|}{2r^{\frac{3}{2}}\epsilon_{Q}}\left(1-\sqrt{1-\frac{4r^{\frac{3}{2}}\epsilon_{Q}}{3\lambda_{1}^{2}}\|\bm{\tilde{d}}\|_{2}}\right),$ (25)
while all trajectories start inside $\bm{\mathrm{B}}_{\bm{y}}(\rho_{+})$ are bounded where
$\rho_{+}=\frac{3|\lambda_{1}|}{2r^{\frac{3}{2}}\epsilon_{Q}}\left(1+\sqrt{1-\frac{4r^{\frac{3}{2}}\epsilon_{Q}}{3\lambda_{1}^{2}}\|\bm{\tilde{d}}\|_{2}}\right).$ (26)

### v3

@@ I Introduction
To determine if a Galerkin model is long-term bounded, Schlegel and Noack [1] developed a “trapping theorem” with necessary and sufficient conditions for long-term model stability for systems that exhibit quadratic, energy-preserving nonlinearities. This theorem can be used as an effective diagnostic for evaluating the stability of reduced-order models. Moreover, the theorem was subsequently incorporated into system identification methods, for producing data-driven models that are provably long-term bounded [2, 14, 15, 16]. The ability to promote global stability guarantees has broad potential for data-driven models, and has already been extended to neural-network-based system identification methods [17]. Recently, Goyal et al.[18] extended the trapping theorem in quadratic systems to generalized energy-preserving nonlinearities, achieved through constraints of a special parametrization that guarantee the stability of the learned model by construction. For general quadratic systems, Kramer [19] discusses the condition for such systems to be locally stable by means of a Lyapunov function and provide better estimates of the radii of the domain of attraction by a novel optimization-based approach. For systems whose dynamics can be described by more general polynomial functions, the sum-of-squares of polynomials techniques provide a constructive method for generating Lyapunov functions to achieve global stability and control of dynamical systems and ROMs [20, 21]. However, a sum-of-squares polynomial Lyapunov function for a given system must be searched for and its existence is not guaranteed.
@@
In this work, we develop a theorem in the context of projection-based ROMs that enables us to promote local stability of data-driven models with quadratic nonlinearities. The theorem provides sufficient condition under which linear-quadratic systems with a relaxed quadratically energy-preserving structure are locally stable. We can also reproduce the Schlegel and Noack theorem [1] by tightening the relaxation of the energy-preserving constraints. Formalizing stability under weak breaking of the energy-preserving nonlinearity is critical because many flows exhibit boundary conditions such as inflow/outflow conditions. Providing conditions under which this class of systems is guaranteed to be stable is helpful for designing fluid flow control.
@@ II Reduced-order modeling
The density $\rho$ and viscosity $\nu$ are assumed constant, and the pressure is denoted $p$. These equations are expected to accurately approximate incompressible flow of Newtonian fluids with constant density, kinematic viscosity, and inflow/outflow boundary conditions. Other complex fluid flows such as ocean currents, viscoelastic fluids, or ferrofluids are described by models that produce higher order polynomial nonlinearities or trigonometric nonlinear terms, but may be amenable to the quadratic analysis here if regimes exist where higher order nonlinearities are inactive or weak or can be transformed into quadratic nonlinearities [31].
The goal of a projection-based ROM is to transform this high-dimensional system into a lower-dimensional system of size $r\ll n$ that captures the essential dynamics. One way to reduce the set of governing equations to a set of ordinary differential equations is by decomposition into a desired low-dimensional orthogonal basis $\{\bm{\chi}_{i}(\bm{x})\}$ in a process commonly referred to as Galerkin expansion:
@@
The usage of Einstein notation is throughout this paper so repeated indices are implicitly summed over. Thus here the sum over $i$ is implied and $i\in\{1,...,r\}$. Here, $\overline{\bm{u}}(\bm{x})$ is a base flow which fulfills stationary boundary conditions, $\bm{\chi}_{i}(\bm{x})$ are spatial modes of the expansion for the fluctuation $\bm{u}^{\prime}(\bm{x},t)=\bm{u}(\bm{x},t)-\overline{\bm{u}}(\bm{x})$, and $a_{i}(t)$ describe how the amplitude of these modes vary in time. The proper orthogonal decomposition (POD) [11, 32] is frequently used to obtain the basis, since the modes $\bm{\chi}_{i}(\bm{x})$ are orthogonal and ordered by maximal energy content. The governing equations are then Galerkin projected onto the basis $\{\bm{\chi}_{i}(\bm{x})\}$ by substituting Eq. (2) into the PDE of Eq. (1) and using inner products with $\bm{\chi}_{k}$ to remove the spatial dependence. Orthogonal projection onto POD modes is the simplest and most common procedure, resulting in POD-Galerkin models, although Petrov-Galerkin projection [5, 6] improves model performance in some cases.
@@ II.1 Stability definitions
We begin with the definition of boundedness. In this paper, we exclusively consider stability in the form of long-term boundedness of dynamical trajectories. Additionally, we will use globally long-term bounded to describe dynamical systems for which all trajectories are long-term bounded. Locally long-term bounded will refer to a system exhibiting dynamical trajectories that are bounded if they are initialized within some compact region of $\mathbb{R}^{N}$.
@@
$\dot{\bm{x}}=\bm{f}(\bm{x})$ (3)
where map $f:\mathcal{D}\rightarrow\mathbb{R}^{n}$ is locally Lipschitz from an open and connected set $\mathcal{D}\subset\mathbb{R}^{n}$ onto $\mathbb{R}^{n}$. If such points exist, the equilibrium points or fixed points of the system (3) are defined as $\bar{\bm{x}}\in\mathcal{D}$ such that $f(\bar{\bm{x}})=0$. For convenience, we will state all definitions and theorems for the case when $\bar{\bm{x}}=0$, the origin of $\mathbb{R}^{n}$. It can be proved that there is no loss of generality in doing so. When discussing the stability of dynamical systems, it is a common practice to use phase space for the intuitive picture of solutions. Now we define a flow on phase space.
Definition 1.
The solutions of Eq. (3) define a flow, $\phi(\bm{x},t)$ or $\phi_{t}(\bm{x})$, such that $\phi_{t}(\bm{x})$ is the solution of this differential equation at time $t$ with initial value $\bm{x}$. Hence,
$\frac{\mathrm{d}}{\mathrm{d}t}\phi(\bm{x},t)=f(\phi(\bm{x},t)),$
for all $t$ such that the solution through $\bm{x}$ exists and $\phi(\bm{x},0)=\bm{x}$.
Note that the solution $x(t)$ with $x(0)=x_{0}$ by this notation is $\phi(x_{0},t)$ or $\phi_{t}(x_{0})$. We then give the definitions of terms that describe the stability of dynamical systems that will be discussed in this paper using these notations.
@@ 1. $\bullet$
    bounded if there exists a constant $\rho>0$ that is independent of the initial time 00, and for every $c\in(0,\rho)$, there is $\mu=\mu(c)>0$ such that
    $\|{\bm{x}}(0)\|\leq c\ \Rightarrow\ \|{\bm{x}}(t)\|\leq\mu,\ \forall t\geq 0,$ (4)
@@ 2. $\bullet$
    globally bounded if (4) holds for arbitrarily large $\rho$.
  3. $\bullet$
    If there exists some $\eta$ such that $\eta\geq\mu(c)$ for every $c\in(0,\rho)$ in the definition above, we can define an upper bound stability region, i.e. a ball center at the origin with radius $\eta$, since
    $\|{\bm{x}}(0)\|\leq c\ \Rightarrow\ \|{\bm{x}}(t)\|\leq\eta,\ \forall t\geq 0,$ (5)
    meaning that there is an upper bound on how far trajectories can ever be from the origin.
  4. $\bullet$
    We will also use the term “domain of boundedness” to denote the set of all $x(0)=x_{0}$ such that the solutions $x(t)$ of a given dynamical system are bounded.
While local or global boundedness is very general, we will show that for the case of quadratically nonlinear systems, boundedness can be addressed by considering a more specialized class of stability via trapping regions. We will use the Lyapunov function to help define the trapping region, so first we give the definition of the Lyapunov function.
@@ 4. $\bullet$
Let $V:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be a continuously differentiable function. Then we can define the derivative of $V(x)$ along trajectories with respect to time [33, 34],
$\dot{V}=\frac{\mathrm{d}V}{\mathrm{d}t}=\bm{f}(\bm{x})\cdot\nabla V.$
Then $V$ is
  1. $\bullet$
    a Lyapunov function for the value of $c\in\mathbb{R}$ if $\dot{V}(\bm{x})\leq 0$ at all points of $V(\bm{x})=c$,
  2. $\bullet$
    a strict Lyapunov function for $c$ if $\dot{V}(\bm{x})<0$ at all points of $V(\bm{x})<c$.
An important case is when there exists some $a,b\in\mathbb{R}$ such that $V$ is Lyapunov or strict Lyapunov for all $c\in[a,b]$ and if so we denote the set $\{\bm{x}\in\mathbb{R}^{n}|V(\bm{x})\leq a\}$ by $\Omega_{V,a}$. Then each trajectory starting inside $\Omega_{V,b}$ either enters the set $\Omega_{V,a}$ or tends to the set given by $\{\bm{x}\in\mathbb{R}^{n}|\dot{V}(\bm{x})=0\}$. We will later see that if $\Omega_{V,a}$ is bounded, it is a trapping region. Now we give the definition of a trapping region.
Definition 4.
Consider the dynamical system of (3) and let $\phi_{t}(\bm{x})$ denote a position of a solution of the system at time $t$ that started at $\bm{x}\in\mathcal{D}$. Then a set $\mathcal{N}\subset\mathcal{D}$ is
@@ 2. $\bullet$
    an attracting trapping region if there exists a compact set $S\subset\mathcal{D}$ and time $T$ such that $\mathcal{N}\subset S$ is a trapping region and
    $\phi_{t}(S\setminus\mathcal{N})\subset\mathrm{int}(\mathcal{N}),\ \forall t>T,$ (7)
@@ 3. $\bullet$
Figure 1: Sketch of an attracting trapping region. $\bm{\Gamma}_{1}$ (in red, solid) and $\bm{\Gamma}_{2}$ (in blue, solid) are two trajectories starting outside an attracting trapping region $\bm{\mathcal{N}}$ where all trajectories will eventually enter in and stay there forever. $V$ is a Lyapunov function and each dashed closed surface $V=V(\bm{x})=c_{i}\ (i=1,2,3,4)$ is a Lyapunov surface with $c_{1}>c_{2}>c_{3}>c_{4}$. For each Lyapunov surface outside $\bm{\mathcal{N}}$, $\bm{\Gamma}_{2}$ only passes through at most once, while $\bm{\Gamma}_{1}$ crosses $V=c_{3}$ more than once before each enters into $\mathcal{\bm{N}}$. If we consider the Lyapunov function as a form of energy, along $\bm{\Gamma}_{2}$, the energy is always decreasing before entering $\bm{\mathcal{N}}$. If all trajectories starting outside $\bm{\mathcal{N}}$ behave the same as $\bm{\Gamma}_{2}$, then $\bm{\mathcal{N}}$ is a monotonically attracting trapping region, otherwise an attracting trapping region.
@@ II.2 Flows with energy-preserving nonlinearities
$E_{i}$, $L_{ij}$, and $Q_{ijk}$ are tensors of time-independent coefficients, obtained from spatial inner products between the $\bm{\chi}_{i}(\bm{x})$ and the spatiotemporal operators that define the model dynamics. We have assumed that the pressure gradient term in Eq. (1) has been removed as in Rowley et al[35]. Moreover, Eq. (9) can be proved that it is locally Lipschitz since it is continuously differentiable over any compact domain in a Euclidean space; therefore the existence and uniqueness of the initial value problem are guaranteed. The class of systems considered for the original trapping theorem are those with energy-preserving nonlinearity, which is defined as
@@
The $Q_{ijk}$ are symmetric in swapping $j$ and $k$ without loss of generality. To see why this condition preserves the energy, consider $\overline{\bm{u}}(\bm{x})$ being the mean flow and we can introduce the standard kinetic energy $K$,
@@ II.3 Schlegel and Noack trapping theorem
The Schlegel and Noack trapping theorem provides necessary and sufficient conditions for energy-preserving, effectively nonlinear, quadratic systems to exhibit a monotonically trapping region, $\bm{B}(\bm{m},R_{m})$, a ball centered at $\bm{m}\in\mathbb{R}^{r}$ with radius $R_{m}$. Outside this region the rate of change of energy $K$ is negative everywhere, producing a Lyapunov function that renders this system globally stable. The trapping theorem begins by recentering the origin by a now arbitrary constant vector $\bm{m}$ so that we can define a shifted energy expressed in terms of the shifted state vector $\bm{y}(t)=\bm{a}(t)-\bm{m}$,
@@
Consider an effectively quadratically nonlinear system. There exists a monotonically trapping ball $\bm{B}(\bm{m},R_{m})$ if and only if the real, symmetric matrix $\bm{A}^{S}$ is negative definite (Hurwitz) with eigenvalues $\lambda_{r}\leq\cdots\leq\lambda_{1}<0$; the radius is then given by $R_{m}=\|\bm{d}\|_{2}/|\lambda_{1}|$. [^1] If a system is long-term bounded (not necessarily exhibiting a monotonically trapping region) and effectively nonlinear, the existence of an $\bm{m}$ producing a negative semidefinite $\bm{A}^{S}$ is still guaranteed.
In practice, the goal is then to find an origin $\bm{m}$ so that the matrix $\bm{A}^{S}$ is Hurwitz, guaranteeing a trapping region and global stability. Moreover, we may want a tight bound on the size of this trapping ball, which can be achieved by varying $\bm{m}$ to reduce $R_{m}$. If $R_{m}\gg 1$, the existence of a trapping ball tells us relatively little about the long-time dynamical trajectories except for the guarantee of boundedness. There is also a caveat for systems that do not exhibit effective nonlinearity; effective nonlinearity means that there is no linear subspace in which a trajectory starts and remains indefinitely, so that the quadratic nonlinearities $Q_{ijk}y_{j}y_{k}$ are never activated.
@@ II.4 Analytical estimates of the local stability of linear-quadratic systems
In this section, we use Lyapunov’s direct method to explore the stability of a linear-quadratic system of the form of Eq. (9). We consider the situation where the nonlinearities are no longer exactly energy-preserving for a linear-quadratic system. However, we assume that the breaking of the energy-preserving nonlinearity is bounded in Frobenius norm by a small amount $\epsilon_{Q}$, i.e.,
$\|Q_{ijk}+Q_{jik}+Q_{kji}\|_{F}\leq\epsilon_{Q},$ (18)
@@
where $\otimes$ denotes the standard Kronecker product. Notice that we previously assume that the $Q_{ijk}$ are symmetric in the last two indices, hence $\bm{H}$ is symmetric, i.e., $\bm{H}(\bm{y}_{1}\otimes\bm{y}_{2})=\bm{H}(\bm{y}_{2}\otimes\bm{y}_{1})$ for arbitrary $\bm{y}_{1},\bm{y}_{2}\in\mathbb{R}^{r}$, which does not change the dynamics [37]. To generalize the discussion beyond the shifted energy, we define a Lyapunov function $v(\bm{y})$ given a positive definite matrix $\bm{P}$,
$v(\bm{y})$ $=\frac{1}{2}\|\bm{y}\|^{2}_{\bm{P}}=\frac{1}{2}\langle\bm{y},\bm{y}\rangle_{\bm{P}}=\frac{1}{2}\bm{y}^{T}\bm{P}\bm{y},$ (24)
@@
$\dot{v}(\bm{y})$ $=\bm{y}^{T}\bm{P}\dot{\bm{y}}.$ (25)
This matrix $P$ is any positive definite matrix that forms the equation [38]
$\bm{A}^{T}\bm{P}+\bm{P}\bm{A}+\bm{M}=0,$
where $A$ is Hurwitz and $\bm{M}$ is positive definite. We call such positive definite matrix $\bm{P}$ the Lyapunov matrix. Importantly, the traditional trapping theorem uses the Lyapunov matrix $\bm{P}=\bm{I}$, the identity, but for two-dimensional fluid flows the enstrophy is an invariant with $\bm{P}\neq\bm{I}$ and it can also be immediately used in the following results. Similar invariants exist in MHD, and more general Lyapunov matrices $\bm{P}$ may exist for certain flows. Section IV.3 will illustrate provably locally-stable, data-driven models of the von Karman vortex street using the enstrophy version of the various trapping theorems described in the present work.
Since the domain of boundedness $\bm{\Omega}_{\bm{m}}$ can be complicated geometrically, we consider closed norm balls $\textbf{B}\subseteq\bm{\Omega}_{\bm{m}}$ as the candidates of regions where the solutions of Eq. (23) are bounded,
$\textbf{B}_{\bm{y}}(R)=\{\bm{y}\in\mathbb{R}^{r}\ |\ \|\bm{y}\|_{\bm{P}}\leq R\},\quad\|\bm{y}\|_{\bm{P}}=\sqrt{\bm{y}^{T}\bm{P}\bm{y}}.$
Note that
$v(\bm{y})=\frac{1}{2}\bm{y}^{T}\bm{P}\bm{y}=\frac{1}{2}\bm{z}^{T}\bm{z},\quad\bm{z}=\bm{P}^{\frac{1}{2}}\bm{y}$ (26)
where $\bm{P}^{\frac{1}{2}}$ and its inverse exist (and are also symmetric positive-definite) because $\bm{P}$ is symmetric positive-definite. It follows that
$\textbf{B}_{\bm{y}}(R)=\textbf{B}_{\bm{z}}(R)=\left\{\bm{z}\in\mathbb{R}^{r}\ |\ \|\bm{z}\|_{2}\leq R\right\}.$ (27)
We now provide a lower-bound for the domain of boundedness by the set $\textbf{B}_{\bm{y}}$. The task then becomes to make this bound as tight as possible, which is, to find $\sup{R}$ as the estimate of the local stability radius. This leads to the following theorem.
@@
    $\bm{\tilde{A}}^{S}=\frac{1}{2}\left(\bm{P}^{\frac{1}{2}}\bm{A}\bm{P}^{-\frac{1}{2}}+\bm{P}^{-\frac{1}{2}}\bm{A}^{T}\bm{P}^{\frac{1}{2}}\right)$ is Hurwitz, i.e. $\lambda_{r}(\bm{\tilde{A}}^{S})\leq...\leq\lambda_{1}(\bm{\tilde{A}}^{S})<0$.
@@
    $\lambda^{2}_{1}-\frac{4\epsilon_{Q}}{3}\|\bm{\tilde{d}}\|_{2}\geq 0$, where $\epsilon_{Q}=\|\bm{\tilde{H}}_{0}\|_{F}$, $\bm{\tilde{H}}_{0}=\bm{P}^{-\frac{1}{2}}\mathcal{T}(\tilde{Q}_{ijk}+\tilde{Q}_{ijk}+\tilde{Q}_{kji})(\bm{P}^{-\frac{1}{2}}\otimes\bm{P}^{-\frac{1}{2}})$, $\bm{\tilde{Q}}=\bm{PQ}$ and $\tilde{\bm{d}}=\bm{P}^{\frac{1}{2}}\bm{d}$.
@@
$\rho_{-}=\frac{3|\lambda_{1}|}{2\epsilon_{Q}}\left(1-\sqrt{1-\frac{4\epsilon_{Q}}{3\lambda_{1}^{2}}\|\bm{\tilde{d}}\|_{2}}\right),$ (28)
while all trajectories starting inside $\bm{\mathrm{B}}_{\bm{y}}(\rho_{+})$ are bounded where
$\rho_{+}=\frac{3|\lambda_{1}|}{2\epsilon_{Q}}\left(1+\sqrt{1-\frac{4\epsilon_{Q}}{3\lambda_{1}^{2}}\|\bm{\tilde{d}}\|_{2}}\right).$ (29)